OpenStudy (anonymous):

work as directed give the equation for the asymptotes of the hyperbola x^2/49 - y^2/12 =1 asymptotes equations are?

OpenStudy (amistre64):

the slope of a line is defined as y/x the slope of the asymptotes are defined by the sqrt of the under y, divided by the under x

OpenStudy (amistre64):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] \[Asymp_{slope}=\frac ba\]

OpenStudy (amistre64):

define the lines by anchoring that +- slope to the h,k center point